mhchen:

Let u be an arbitrary vector from vector space V Prove that -u = (-1)u

mhchen:

List of axioms: Let u, v be vectors k,m be scalars 1. If u and v are in V, then u+v is in V 2. u + v = v + u 3. u + (v + w) = (u+v) + w 4. 0 + u = u + 0 = u 5. u + (-u) = (-u) +u = 0 6. If k is any scalar and u is in V, the ku is in V 7. k(u+v) = ku+kv 8.(k+m)u = ku + mu 9. k(mu) = (km)u 10. 1*u = u

mhchen:

Axiom 5 is the only one that has a (-u) in it, but I'm not sure how to derive (-1)u=-u from it.

Narad:

What do you think about 6?

mhchen:

I could use 6 to show that (-1)u is in V. But I can't say that (-1)u=-u (like seriously what the heck math makes no sense at this point)